A satellite is launched to orbit the Earth at an a
Answer #1

b)

Speed of satellite = (GM/R )1/2

R = r + a

r = radius

a = distance of satellite= 2.85*10^7   

V2 = (6.67*10-11 *5.97*1024)/(2.85*107+ 6.38*106) = 11416255.73

V = 3378.795 m/s